# Intelligent Metasurface-Enabled Integrated Sensing and Communication: Unified Framework and Key Technologies

### v1

@@ II Unified Framework for IM-enabled ISAC
Baseband digital precoding Yes Yes No No
@@
Signal from feed source Info-bearing Info-bearing Carrier only Info-bearing
@@
Element reconfiguration Amplitude/Phase Amplitude only Phase Phase
Element constraint Lorentzian Amplitude discrete Unit-modulus Unit-modulus
@@ II-A Existing IM-enabled ISAC Transceiver Design
Stacked Intelligent Metasurface: SIM utilizes multiple cascaded transmissive metasurface layers that facilitate sophisticated wave-domain signal processing through their enhanced EM wave manipulation capabilities. As shown in Fig. 1b, SIM mainly comprises multiple parallel metasurfaces with sub-wavelength separation, fed by external RF chains [8]. When signals from the RF chains propagate through these metasurfaces, a cascaded coupling effect is established. This effect begins with the initial signals from the RF chains fully engaging all elements on the first layer. Subsequently, elements on each metasurface re-radiate EM waves, which in turn serve as the input illuminating all elements of the next metasurface in the cascade. Wave-domain signal processing thus occurs sequentially as these waves traverse the entire multi-layer structure, with phase shifts at each meta-element on each layer reconfiguring the wavefront, akin to a physical neural network. This multi-layer structure offers vastly increased spatial DoF, enabling the generation of highly complex and tailored beam patterns for enhanced communication and sensing performance [7]. On the other hand, the intricate interlayer coupling inherent in SIM presents considerable challenges for ISAC beamforming, necessitating meticulous optimization of all elemental phase shifts to achieve desired dual-functionality, which will be elaborated on in Section III-C.
@@
Dynamic Metasurface Antenna: DMA is an emerging concept within the paradigm of “radiative metasurfaces”, which uniquely performs sophisticated analog signal processing directly within a waveguide-integrated radiating structure [9]. As shown in Fig. 1c, the DMA architecture consists of multiple configurable meta-elements situated on a set of one-dimensional waveguides (typically implemented as microstrip lines). Each RF chain is connected to a distinct waveguide and exclusively controls the meta-elements situated on that path, forming a partially connected hybrid structure similar to conventional array antennas [9]. Under this distinctive architecture, the independent signal streams from RF chains are fed in parallel to their respective waveguides. Each stream sequentially traverses a series of meta-elements, where each element reconfigures and radiates the incoming signal into the wireless channel based on its frequency response. The signal manipulation in the DMA is predominantly influenced by two factors: the propagation characteristics within the waveguide and the reconfigurable response of the radiating elements [9]. The former determines the propagation attenuation and phase shifts, which originate from the dielectric loss and conductor loss that depend on the material properties of the microstrip line and carrier frequency. The latter pertains to the element’s reconfigurable state, which directly dictates its radiation amplitude and phase. This state is dynamically tunable; however, the achievable amplitude and phase are typically interdependent, generally adhering to Lorentzian constraints [10]. Nevertheless, DMA can synthesize highly complex radiation patterns through precise, collective control of these sub-wavelength spacing tunable elements, enabling advanced generation of tailored concurrent communication and sensing beams.
@@
Reconfigurable Holographic Surface: RHS is an IM-enabled transceiver that incorporates the philosophies from leaky-wave and holographic antennas [11]. As illustrated in Fig. 1d, RHS mainly comprises a parallel-plate waveguide equipped with sub-wavelength amplitude-controlled meta-elements. Similar to DMA, RF signals are fed into the parallel-plate waveguide and propagate sequentially past each radiation element. Therein the guided surface waves progressively convert into leaky waves, radiating into the wireless channel. Unlike Holographic MIMO (HMIMO), which typically involves (quasi-)continuous apertures (e.g., as in [13]) and is further detailed in Section III-A, the “holographic” aspect of RHS is rooted in optical holographic principles for wavefront reconstruction. In this holographic paradigm, the guided wave traveling along the waveguide serves as the reference wave, while the desired beam pattern or EM field distribution acts as the object wave. The configurable amplitude profile of the RHS elements then functions as a holographic pattern, which records the interference produced by the coherent superposition between these two waves. When illuminated by the reference wave, this holographic pattern reconstructs the corresponding object wave, thereby achieving the required precoding/combining functionality. Therefore, by modulating the voltage applied to each element to meticulously regulate its radiation amplitude, RHS can dynamically reconfigure its holographic pattern. This allows it to generate required object waves (beams) for both communication and sensing. Compared to traditional array-based architecture, RHS can provide lower power consumption and hardware cost by eliminating the need for phase-shifting circuit [11].
@@ II-B Unified Framework for IM-enabled ISAC Systems
Multiple RF Chains Digital Precoding Matrix $\blacksquare$ $\blacktriangle$
@@
As illustrated in Figs. 2a and 2b, the first slice in our unified framework is the baseband processing, where the Baseband Processing Matrix ($\mathbf{V}$) processes the input data streams destined for the RF chains. Within our unified framework, this process is categorized into three primary types according to the presence of digital precoding and the number of RF chains in the IM-enabled ISAC system, as clarified in Table II. For architectures equipped with a digital precoder, such as DMA and RHS, the baseband processing maps multiple data streams to the corresponding RF chains to enable spatial multiplexing. For other architectures that forgo digital precoding, the baseband process is concerned with power management. In the case of a multi-RF-chain system like SIM, this is represented by a power allocation vector. In a single-RF-chain configuration such as RIS, baseband processing is reduced to simply assigning a scalar power level to the carrier signal.
@@
The final slice in our framework is the reconfiguration process, where the metasurfaces alter the wavefront by applying element-wise phase and/or amplitude modifications (see Figs. 2e and 2f). This is characterized by the Reconfiguration Matrix ($\mathbf{Q}_{l}$). In multi-layer architectures like SIM, the signal processing takes on a cascaded form, where the RF feeding matrix $\mathbf{T}_{l}$ and the reconfiguration matrix $\mathbf{Q}_{l}$ for each layer act as a pair, representing the sequential processes of signal transmission to, and reconfiguration at, the $l$-th layer. When inter-element coupling effects are neglected, the reconfiguration matrix is typically diagonal. Each diagonal entry thus represents a specific complex coefficient applied by a single meta-element to its local incident signal. The ability to adjust these complex coefficients is the key enabler of reconfigurability in an IM-enabled ISAC system. As summarized in Table II, this tuning capability manifests as distinct EM control modalities, each regulated by a specific constraint. These modalities include amplitude-only control within a defined range or set (e.g., RHS), phase-only control with a unit-modulus constraint (e.g., RIS and SIM), and a coupled amplitude-and-phase response governed by a Lorentzian constraint (e.g., DMA).
@@ III-C IM-enabled ISAC Beamforming
For RHS-enabled ISAC systems, the beamforming design balances beam pattern matching for sensing and communication with SINR threshold constraints by optimizing baseband digital precoders and the RHS reconfiguration matrix. Decomposing the optimization into subproblems (e.g., digital precoding and reconfiguration) and employing techniques such as semi-definite relaxation and alternating optimization effectively handles the coupling between the propagation matrix and the amplitude-constrained reconfiguration matrix. [11].
@@ IV Conclusions
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (NSFC) under grant number 62401054, in part by the Open Project Program of State Key Laboratory of CNS/ATM under Grant number 2024B12, and in part by Young Elite Scientists Sponsorship Program by CAST under Grant number 2022QNRC001.

### v2

@@ II Unified Framework for IM-enabled ISAC
Baseband digital precoding Yes Yes No Optional (typically no)
@@
Signal from feed source Information-bearing Information-bearing Carrier only Information-bearing
@@
Element reconfiguration Coupled amplitude-phase Amplitude only Phase only Phase only
Element constraint Lorentzian constraint Amplitude discrete Unit-modulus Unit-modulus
@@ II-A Existing IM-enabled ISAC Transceiver Design
Stacked Intelligent Metasurface: SIM utilizes multiple cascaded transmissive metasurface layers that facilitate sophisticated wave-domain signal processing through their enhanced EM wave manipulation capabilities. As shown in Fig. 1b, SIM mainly comprises multiple parallel metasurfaces with sub-wavelength separation, fed by external RF chains [7]. When signals from the RF chains propagate through these metasurfaces, a cascaded coupling effect is established. This effect begins with the initial signals from the RF chains fully engaging all elements on the first layer. Subsequently, elements on each metasurface re-radiate EM waves, which in turn serve as the input illuminating all elements of the next metasurface in the cascade. Wave-domain signal processing thus occurs sequentially as these waves traverse the entire multi-layer structure, with phase shifts at each meta-element on each layer reconfiguring the wavefront, akin to a physical neural network. This multi-layer structure offers vastly increased spatial DoF, enabling the generation of highly complex and tailored beam patterns for enhanced communication and sensing performance [8]. On the other hand, the intricate interlayer coupling inherent in SIM presents considerable challenges for ISAC beamforming, necessitating meticulous optimization of all elemental phase shifts to achieve desired dual-functionality, which will be elaborated on in Section III-C. The SIM architecture considered in this work follows the widely adopted baseline implementation without baseband digital precoding. Nevertheless, an optional variant incorporating a digital precoder upstream of the RF chains can also be envisioned, whose modeling is accommodated within the proposed unified framework and elaborated in the subsequent subsection.
@@
Dynamic Metasurface Antenna: DMA is an emerging concept within the paradigm of “radiative metasurfaces”, which uniquely performs sophisticated analog signal processing directly within a waveguide-integrated radiating structure [9]. As shown in Fig. 1c, the DMA architecture consists of multiple configurable meta-elements situated on a set of one-dimensional waveguides (typically implemented as microstrip lines). Each RF chain is connected to a distinct waveguide and exclusively controls the meta-elements situated on that path, forming a partially connected hybrid structure similar to conventional array antennas [10]. Under this distinctive architecture, the independent signal streams from RF chains are fed in parallel to their respective waveguides. Each stream sequentially traverses a series of meta-elements, where each element reconfigures and radiates the incoming signal into the wireless channel based on its frequency response. The signal manipulation in the DMA is predominantly influenced by two factors: the propagation characteristics within the waveguide and the reconfigurable response of the radiating elements [9]. The former determines the propagation attenuation and phase shifts, which originate from the dielectric loss and conductor loss that depend on the material properties of the microstrip line and carrier frequency. The latter pertains to the meta-elements’ resonant Lorentzian frequency response. Under a normalized narrowband surrogate around the operating frequency, it reduces to a frequency-independent complex-plane parameterization where the radiated amplitude is intrinsically coupled to the desired phase shift [9]. Nevertheless, DMA can synthesize highly complex radiation patterns through precise, collective control of these sub-wavelength spacing tunable elements, enabling advanced generation of tailored concurrent communication and sensing beams.
@@
Reconfigurable Holographic Surface: RHS is an IM-enabled transceiver that incorporates the philosophies from leaky-wave and holographic antennas [11]. Structurally, an RHS shares similarities with a DMA: as illustrated in Fig. 1d, it comprises a parallel-plate waveguide where RF signals are fed and propagate sequentially past each radiation element. The key distinction lies in the function of the sub-wavelength meta-elements: instead of resonance tuning as in DMA, RHS primarily controls the local leakage rate (i.e., radiated amplitude) along the waveguide. Unlike Holographic MIMO (HMIMO), which typically involves (quasi-)continuous apertures (e.g., as in [14]) and is further detailed in Section III-A, the “holographic” aspect of RHS is rooted in optical holographic principles for wavefront reconstruction. In this holographic paradigm, the guided wave traveling along the waveguide serves as the reference wave, while the desired beam pattern or EM field distribution acts as the object wave. The configurable amplitude profile of the RHS elements then functions as a holographic pattern, which records the interference produced by the coherent superposition between these two waves. When illuminated by the reference wave, this holographic pattern reconstructs the corresponding object wave, thereby achieving the required precoding/combining functionality. Therefore, by modulating the voltage applied to each element to meticulously regulate its radiation amplitude, RHS can dynamically reconfigure its holographic pattern. This allows it to generate required object waves (beams) for both communication and sensing. Compared to traditional array-based architecture, RHS can provide lower power consumption and hardware cost by eliminating the need for phase-shifting circuit [11].
@@ II-B Unified Framework for IM-enabled ISAC Systems
Multiple RF Chains Digital Precoding Matrix $\blacksquare$ $\blacktriangle$ $(\bigstar)*$
@@
As illustrated in Figs. 2a and 2b, the first slice in our unified framework is the baseband processing, where the Baseband Processing Matrix ($\mathbf{V}$) processes the input data streams destined for the RF chains. Within our unified framework, this process is categorized into three primary types according to the presence of digital precoding and the number of RF chains in the IM-enabled ISAC system, as clarified in Table II. For architectures equipped with a digital precoder, such as DMA and RHS, the baseband processing maps multiple data streams to the corresponding RF chains to enable spatial multiplexing. For other architectures that forgo digital precoding, the baseband process is concerned with power management. In a single-RF-chain configuration such as RIS, baseband processing reduces to simply assigning a scalar power level to the carrier signal. For the optional (typically not adopted) SIM variant with baseband digital precoding, the baseband processing becomes a digital precoding matrix applied upstream of the RF chains.
@@
The final slice in our framework is the reconfiguration process, where the metasurfaces alter the wavefront by applying element-wise phase and/or amplitude modifications (see Figs. 2e and 2f). This is characterized by the Reconfiguration Matrix ($\mathbf{Q}_{l}$). In multi-layer architectures like SIM, the signal processing takes on a cascaded form, where the RF feeding matrix $\mathbf{T}_{l}$ and the reconfiguration matrix $\mathbf{Q}_{l}$ for each layer act as a pair, representing the sequential processes of signal transmission to, and reconfiguration at, the $l$-th layer. When inter-element coupling effects are neglected, the reconfiguration matrix is typically diagonal. Each diagonal entry thus represents a specific complex coefficient applied by a single meta-element to its local incident signal. The ability to adjust these complex coefficients is the key enabler of reconfigurability in an IM-enabled ISAC system. As summarized in Table II, this tuning capability manifests as distinct EM control modalities, each regulated by a specific constraint. These modalities include amplitude-only control within a prescribed range (e.g., RHS), phase-only control under a unit-modulus constraint (e.g., RIS and SIM), and coupled amplitude–phase control governed by a Lorentzian constraint (e.g., DMA).
@@
Moreover, the framework can be extended to time‑varying IMs by modeling the reconfiguration as a time‑periodic operator and expanding it into Fourier harmonics, after which (1)-(2) are applied to each harmonic without changing their form. This provides a unified description of spatial beamforming and spectral shaping at each generated frequency, consistent with established practices in time-modulated arrays [5, 12].
@@ III-C IM-enabled ISAC Beamforming
For RHS-enabled ISAC systems, the beamforming design balances beam pattern matching for sensing and communication with SINR threshold constraints by optimizing baseband digital precoders and the RHS reconfiguration matrix. Decomposing the optimization into subproblems (e.g., digital precoding and reconfiguration) and employing techniques such as semi-definite relaxation and alternating optimization effectively handles the coupling between the propagation matrix and the amplitude-constrained reconfiguration matrix [11].
In contrast to the amplitude-only control of RHS, DMA-enabled ISAC beamforming must account for a Lorentzian hardware constraint that intrinsically couples amplitude and phase at each meta-element, thereby posing a distinct challenge for joint ISAC design. A practical remedy is to handle the Lorentzian constraint via an equivalent unit‑modulus proxy variable, which recasts the design as a standard constant‑modulus optimization [9, 10]. In this formulation, the feasible set of the optimization variable becomes the product of complex unit circles, which is a Riemannian submanifold. The resulting problem can be effectively handled by manifold-optimization methods such as the Riemannian Conjugate Gradient (RCG).
@@ IV Conclusions
Biographies
Shunyu Li received the B.S. degree from the School of Information and Electronics, Beijing Institute of Technology (BIT), Beijing, China, in 2021, and the M.Sc. degree from the Faculty of Science and Engineering, The University of Manchester, U.K., in 2022. He is currently working toward the Ph.D. degree with the School of Information and Electronics, BIT.
Tianqi Mao (Member, IEEE) is currently an Associate Professor with the School of Interdisciplinary Science, Beijing Institute of Technology, Beijing, China. His research interests include modulation, waveform design and signal processing for wireless communications, integrated sensing and communication, terahertz communications, and visible light communications.
Guangyao Liu received the B.S. degree in electronics information engineering from Dalian University of Technology, Dalian, China, in 2019. He is currently working toward the Ph.D. degree with the School of Electronic and Information Engineering, Beihang University.
Fan Zhang received the B.S. (Hons.) degree from in 2022 the Department of Electronic Engineering, Tsinghua University, Beijing, China, where she is currently working toward the Ph.D. degree. Her main research interests include wireless communications, signal processing, waveform design in joint sensing, and communication systems.
Ruiqi (Richie) Liu (Senior Member, IEEE) is a master researcher in the wireless and computing research institute of ZTE Corporation. He is deeply involved in specifying 5G standards through 3GPP and 6G framework through ITU, where he served as a rapporteur or chair. He is a Voting Member of the IEEE ComSoc Industry Communities Board.
Meng Hua (Member, IEEE) is a Research Associate with the Department of Electrical and Electronic Engineering, Imperial College London, United Kingdom.
Zhen Gao (Member, IEEE) received the Ph.D. degree in communication and signal processing from Tsinghua National Laboratory for Information Science and Technology, Department of Electronic Engineering, Tsinghua University, China, in 2016. He is currently a Professor with the School of Interdisciplinary Science, Beijing Institute of Technology, Beijing.
Qingqing Wu (Senior Member, IEEE) is currently an Associate Professor with the Department of Electronic Engineering, Shanghai Jiao Tong University, China. He was listed as the Clarivate ESl Highly Cited Researcher in 2021 and 2022, the Most Influential Scholar Award in Al-2000 by Aminer in 2021, and World’s Top 2 percent Scientist by Stanford University in 2020 and 2021.
George K. Karagiannidis (Fellow, IEEE) is currently a Professor with the ECE Department of Aristotle University of Thessaloniki, Greece, and the Head and the Founder of the Wireless Communications & Information Processing (WCIP) Group. He is one of the highly-cited authors across all areas of Electrical Engineering, recognized from Clarivate Analytics as a Highly-Cited Researcher in the ten consecutive years 2015–2024.